if f(x)=x^2+3 and g(x)=3x-1
then find the following

1a. (f+g)(x)

1b. (f+g)(3)

1c. (f-g)(x)
1
Expert's answer
2012-12-19T10:23:22-0500
1a
(f+g)(x)=f(x)+g(x)=x^2+3+3x-1=x^2+3x+2

1.b
(f+g)(x)=x^2+3x+2
(f+g)(3)=3^2+3*3+2=9+9+2=20

1c.
(f-g)(x)= x^2+3-(3x-1)=x^2-3x+4